# Spontaneous Decoherence from Imaginary–Order Spectral Deformations

### v2

@@ 1 Introduction
with eigenstate amplitudes $e^{-itE^{1+i\beta}}$. The exponent introduces phases depending on $\beta\log E$, and the resulting oscillatory structure suppresses interference between distinct energies. The mechanism is deterministic and internal to the system: it does not rely on baths, measurements, or nonlinearity. The Hilbert–space inner product, projection operators, and probability assignments via the Born rule remain exactly as in standard quantum mechanics. Only the dynamical phases are modified.
@@
The present proposal is intended as a phenomenological mechanism for spontaneous decoherence within strictly unitary quantum dynamics, encoded by the imaginary–order spectral deformation $H^{1+i\beta}$. It does not aim to solve the measurement problem, to replace standard environment–induced decoherence, or to derive $\beta$ from a specific microscopic or quantum–gravity model. Rather, it provides a compact way to parameterize logarithmic spectral corrections and to assess their potential impact on coherence in realistic platforms.
@@ 6 Illustrative Examples
We now illustrate how the deformation $H^{1+i\beta}$ acts on several simple Hamiltonians. The goal is not exhaustive analysis, but to show concretely how spectral structure feeds into decoherence rates.
@@ 7.1 General scaling and decoherence envelopes
with corresponding timescale $\tau_{\mathrm{dec}}\sim 1/\Gamma_{mn}$. If environmental decoherence produces a standard coherence envelope $C_{\mathrm{std}}(t)$, the present mechanism can be viewed as supplying an additional deterministic decay envelope, so that
@@
at least over timescales where the non–stationary–phase estimate is valid. In an experiment, one would first calibrate $C_{\mathrm{std}}(t)$ using environmental models and then fit the residual decay to extract or bound $\Gamma_{mn}$, and hence $\beta$. This logic parallels the way collapse–model parameters are bounded by precision experiments [6].
@@ 8.2 Environmental / open–system models
The following table highlights key structural differences between the present mechanism and existing approaches:
@@ Acknowledgments
The author thanks colleagues in quantum foundations and quantum gravity whose discussions helped refine the presentation of imaginary–order spectral deformations and their relation to existing decoherence models.

### v3

@@ 1 Introduction
with eigenstate amplitudes $e^{-itE^{1+i\beta}}$. The exponent introduces phases depending on $\beta\log E$, and the resulting oscillatory structure suppresses interference between distinct energies. The mechanism is deterministic and internal to the system: it does not rely on baths, measurements, or nonlinearity. The Hilbert–space inner product, projection operators, and probability assignments via the Born rule remain exactly as in standard quantum mechanics. Unlike operational or geometric approaches in which decoherence arises from kinematic restrictions on accessible observables, the present mechanism modifies the generator of time evolution itself, even though its observable consequences emerge only after spectral averaging or finite experimental resolution. Only the dynamical phases are modified.
@@
The present proposal is intended as a phenomenological mechanism for spontaneous decoherence within strictly unitary quantum dynamics, encoded by the imaginary–order spectral deformation (IOSD) $H^{1+i\beta}$. IOSD does not solve the measurement problem (see Appendix A for a possible interpretive note), to replace standard environment–induced decoherence, or to derive $\beta$ from a specific microscopic or quantum–gravity model. Rather, it provides a compact way to parameterize logarithmic spectral corrections and to assess their potential impact on coherence in realistic platforms.
@@ 6 Illustrative Examples
We now illustrate how the deformation $H^{1+i\beta}$ acts on several simple Hamiltonians. The goal is not exhaustive analysis, but to show concretely how spectral structure feeds into decoherence rates. These examples are not intended to model realistic experimental systems, but to illustrate how different spectral structures feed into interference suppression under the same imaginary–order deformation.
@@ 7.1 General scaling and decoherence envelopes
with corresponding timescale $\tau_{\mathrm{dec}}\sim 1/\Gamma_{mn}$. If environmental decoherence produces a standard coherence envelope $C_{\mathrm{std}}(t)$, the present mechanism may be viewed as suggesting an additional deterministic envelope that could, in principle, be constrained once dominant environmental contributions are independently modeled, so that
@@
at least over timescales where the non–stationary–phase estimate is valid. We emphasize that the exponential envelope in Eq. (38) is not a fundamental loss of norm or nonunitary damping: it is an effective description of interference suppression after spectral averaging and finite experimental resolution, while the underlying evolution generated by $U_{\beta}(t)$ remains exactly unitary. In an experiment, one would first calibrate $C_{\mathrm{std}}(t)$ using environmental models and then fit the residual decay to extract or bound $\Gamma_{mn}$, and hence $\beta$. This logic parallels the way collapse–model parameters are bounded by precision experiments [6].
@@ 8.2 Environmental / open–system models
Table 4 highlights key structural differences between the present mechanism and existing approaches.
@@ Acknowledgments
The author thanks colleagues in quantum foundations and quantum gravity whose discussions and suggestions helped refine the presentation of imaginary–order spectral deformations, their relation to existing decoherence models and develop an interpretive note.
Appendix A Interpretive Note on IOSD
The imaginary–order spectral deformation $H^{1+i\beta}$ leaves the Hilbert–space kinematics of quantum mechanics unchanged, including the inner product, projection operators, and the Born rule, while modifying only the dynamical phases of energy–eigencomponents through factors of the form
$E^{1+i\beta}=E\,E^{i\beta}=E\,e^{i\beta\log E}.$
[…]
Intrinsic Einselection Under IOSD
Let $H$ be a positive self–adjoint Hamiltonian with spectral resolution
$H=\int_{0}^{\infty}E\,d\Pi(E),$
and consider the IOSD evolution
$U_{\beta}(t)=\exp\!\bigl(-itH^{1+i\beta}\bigr)=\int_{0}^{\infty}e^{-itE^{1+i\beta}}d\Pi(E),\qquad\beta\in\mathbb{R}.$
For any coarse–grained observable $A$ with spectral projections $\{P_{a}\}$, the corresponding Heisenberg operators $P_{a}(t)=U_{\beta}(t)^{\dagger}P_{a}U_{\beta}(t)$ generate histories that can be analyzed in the standard decoherent–histories framework. Whenever the relevant components of the initial state $\rho_{0}$ have support on an energy window without stationary points of the IOSD phase $tE^{1+i\beta}$, stationary–phase estimates imply that cross–terms between distinct coarse–grained histories decay in time, so that these histories form an effectively decoherent family in the usual sense. In this regime, pointer states are those whose energy support is sufficiently narrow that the additional phases $E^{i\beta}$ remain approximately coherent over the experimental timescales of interest; superpositions of components with significantly different values of $E\log E$ are dynamically driven into effective superselection sectors by IOSD–induced dephasing. Thus IOSD implements an intrinsic, Hamiltonian–level version of einselection in which robustness is defined spectrally rather than by environmental monitoring.
Optional Single–History Postulate
By design, IOSD does not alter the Born rule or the projection–valued observable structure: probabilities of measurement outcomes are still computed as $\mathrm{Tr}(\rho P)$ with the undeformed inner product. Consequently, IOSD alone—like standard environment–induced decoherence—explains the suppression of interference between branches but does not, by itself, single out one branch as “actual” in an individual run.
If one wishes to supplement IOSD with a single–outcome ontology, a minimal and conventional option is to adopt the standard decoherent–histories move: for any IOSD–decoherent family of coarse–grained histories $\{\alpha\}$, defined by class operators $C_{\alpha}$ constructed from the IOSD evolution, assign probabilities
$p(\alpha)=\mathrm{Tr}\!\bigl(C_{\alpha}\rho_{0}C_{\alpha}^{\dagger}\bigr),$
and postulate that in each individual realization exactly one history $\alpha$ occurs, with long–run frequencies given by $\{p(\alpha)\}$. This postulate does not modify the IOSD dynamics or introduce additional stochastic terms; it simply adds the usual “one actual history” clause familiar from decoherent–histories interpretations. Within this perspective, IOSD furnishes an intrinsic einselection mechanism, while the single–history clause remains an optional interpretive supplement external to the phenomenological content developed in the main text.